The sum of two numbers is 26. The larger number is six more than four times the smaller number. what are the numbers?
Answer by stanbon(75887)

The sum of two numbers is 26. The larger number is six more than four times the smaller number. what are the numbers?
----------------------
Equations:
x + y = 26
x = 4y+6
-----
Substitute:
4y+6 + y = 26
----
5y = 20
----
y = 4
-------------
x = 26-4 = 22
------------------
Cheers,
Stan H.